OpenStudy (anonymous):

Which ordered pair is a solution for this system of inequalities? 2x + 3y > 1 and y < 5/4 x - 1 A.(3, 9) B.(-4, 2) C.(8, -1) D.(-2, -5)

jimthompson5910 (jim_thompson5910):

If A was the answer (that's an if), then (3,9) should satisfy both inequalities

jimthompson5910 (jim_thompson5910):

2x + 3y > 1 2*3 + 3*9 > 1 6+27 > 1 33 > 1 So the first one works, let's try the second

jimthompson5910 (jim_thompson5910):

y < 5/4 x - 1 9 < (5/4)(3) - 1 9 < 15/4 - 1 9 < 3.75 - 1 9 < 2.75 but the second one does NOT work

jimthompson5910 (jim_thompson5910):

So choice A is NOT the answer

jimthompson5910 (jim_thompson5910):

Do the same with the other answer choices to see if the ordered pair works with BOTH inequalities

OpenStudy (anonymous):

c

OpenStudy (anonymous):

okay thank you, Jim.

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

let me know which answer you get

OpenStudy (anonymous):

i got C for my answer.

jimthompson5910 (jim_thompson5910):

Let's see if that's true or not 2x + 3y > 1 2(8) + 3(-1) > 1 16 - 3 > 1 13 > 1 ... True --------------- y < 5/4 x - 1 -1 < (5/4)(8) - 1 -1 < 40/4 - 1 -1 < 10 - 1 -1 < 9 ... also True Since both inequalities are true, (8,-1) is indeed a solution

jimthompson5910 (jim_thompson5910):

so you are correct

OpenStudy (anonymous):

thanks again

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

what about, Solve the system of equations using any method. Find the value of y. 6x + 3y = 12 and 5x + 3y = 0

OpenStudy (anonymous):

options: -20 -12 20 12

jimthompson5910 (jim_thompson5910):

notice we have 3y in both equations

jimthompson5910 (jim_thompson5910):

so if you subtract equation 2 from equation 1, what do you get?

jimthompson5910 (jim_thompson5910):

6x + 3y = 12 5x + 3y = 0 ----------- ??? ??? ???

OpenStudy (anonymous):

1x=12 ?

jimthompson5910 (jim_thompson5910):

good, so x = 12

jimthompson5910 (jim_thompson5910):

now use this to find y 6x + 3y = 12 6(12) + 3y = 12 72 + 3y = 12 I'll let you finish

OpenStudy (anonymous):

y=-20

jimthompson5910 (jim_thompson5910):

good, you got it

OpenStudy (anonymous):

can i message you with more problems? i am taking my final exam and your help is really good

jimthompson5910 (jim_thompson5910):

sure, go ahead